Q3. Assume that T has t-student distribution with v degree of freedom. (a) Obtain P(T> 3.833), with v 8. (b obtain P(0.559 < T< 6.869), with 5 (c)Obtain to.025, with 6. (d Obtain -to.9, with 9 (e) Obtain ta, such that P(0.876 < T< ta)0.1 with 0.025 , Wi 11
Answer #1

from t distribution table

a) T =0.0025

b)P(0.559<T<6.869)=P(T<6.869)-P(T<0.559)=0.9995-0.7000=0.2995

c)t0.025 =2.447

d)-t0.9=-1.383

e)

P(0.876<T<t)=0.1

P(T<t)-P(T<0.876)=0.1

P(T<t)-0.8000=0.1

P(T<t)=0.9

t=1.363